Find the derivative f 0 m(x) of the following function with respect to x:

fm(x) = (mΣ n=1 n^x · x^n )^2
1
Expert's answer
2020-11-16T08:32:31-0500
"f_m(x)=\big(\displaystyle\sum_{ n=1}^m(n^x\cdot x^n)\big)^2"

"f_m'(x)=\bigg(\big(\displaystyle\sum_{n=1}^m(n^x\cdot x^n)\big)^2\bigg)'="

"=2\cdot\displaystyle\sum_{n=1}^m(n^x\cdot x^n)\cdot\displaystyle\sum_{n=1}^m\big(\ln(n)\cdot n^x\cdot x^n+n^{x+1}\cdot x^{n-1}\big)"